OpenStudy (anonymous):

My new TV has a perimeter of 164 inches. Its width is 8 inches less than twice the height. How many inches wide is the TV?

OpenStudy (mathstudent55):

The perimeter is the sum of the lengths of all sides. A rectangle has two equal lengths and two equal widths. Your problem is calling the length width, and the width height, so let's use those terms. P = W + W + H + H or P = 2W + 2H We know the perimeter is 174, so we can write 2W + 2H = 164 Now we can use the information about the width and height to find another equation. "Its width is 8 inches less than twice the height." W = 2H - 8 Now you have a system of two equations. 2W + 2H = 164 W = 2H - 8 Solve the system of equations to find W.

OpenStudy (campbell_st):

here is a simple method the sum of the length and width is 82 since the perimeter of a rectangle is P = 2(h + w) so you can say 82 = h + w let the height = h so the width is 2h - 8 then 82 = h + 2h - 8 or 82 = 3h - 8 solve for h, and you'll have the height... and can then calculate the width...

OpenStudy (anonymous):

h=30

OpenStudy (anonymous):

is the width 52?

OpenStudy (mathstudent55):

Since all you are asked for is the width, you can solve the first equation for h, then substitute it in the second equation. 2W + 2H = 164 W = 2H - 8 Solve the first equation for H. Divide both sides by 2: W + H = 82 Subtract W from both sides: H = 82 - W Now substitute 82 - W for H in the second equation. W = 2H - 8 W = 2(82 - W) - 8 Now solve for W. W = 164 - 2W - 8

OpenStudy (mathstudent55):

Correct.